# Search for Quadruplet Scalars using Boosted Decision Trees at the LHC

### v1

@@ 3.3 Existing Collider Searches
In the model under inspection, the scalars decay either to the gauge bosons directly or to the quintuplet fermions which subsequently decay to gauge bosons, as we have discussed in the previous section. Before discussing possible signatures at the LHC, let us discuss the existing limits from collider experiments on charged scalars. The direct search limit on charged Higgs comes from LEP which is around 80 GeV [ALEPH:2013htx]. Both ATLAS and CMS have conducted extensive searches targeting multiple production and decay modes of the charged scalars. A recent summary of these searches is provided in [Tao:2021tzs, ATLAS:2024itc, Horii:2024qiw, Banerjee:2025qkb]. Most LHC searches target single production of the charged scalars via vector-boson fusion and have put limits in cases where the scalar decays to gauge bosons[ATLAS:2022zuc, ATLAS:2023sua, CMS:2021wlt, ATLAS:2018ceg]. In this case, a lower limit of $\sim 200$ GeV is placed on the masses of charged scalars. Limits also exist on production cross-sections of charged scalars when they directly decay to SM fermions [CMS:2019bfg, ATLAS:2012nhc, ATLAS:2018gfm, ATLAS:2018ntn, ATLAS:2015nkq, ATLAS:2022pbd] or specifically, to quarks [ATLAS:2013uxj]. These searches have placed a lower limit of $\sim 400$ GeV on the charged scalar masses.
Ref. [ATLAS:2021jol] has studied the doubly charged scalar pair production as well as associated production of singly and doubly charged scalars, with each scalar decaying to gauge bosons and limits are placed from the searches in 4 leptons, 3 leptons + $\geq$ 2 jets and 2 leptons $\geq$ 3 jets final states. As no excess has been observed, limits have been placed on the cross-section $\times$ BR (See Fig. 9 of [ATLAS:2021jol]).
@@ 3.4 Fermiophobic and Fermiophilic Signatures
Let us consider the fermiophobic scenario when $M_{\Phi}>M_{\Sigma}$ and the coupling of the scalars with the quintuplet fermion is small, $y_{\nu}<0.005$. In this situation, the branching ratio of the scalars to the SM fermions are negligible. This situation gives rise to the final states with gauge bosons,
$pp~\rightarrow~\phi_{1}^{+}\phi_{1}^{-}$ $\rightarrow$ $W^{+}Z~W^{-}Z$ (20)
$pp\rightarrow\phi^{++}\phi^{--}$ $\rightarrow$ $W^{+}W^{+}W^{-}W^{-}$ (21)
@@
We present the possible final states and the values of cross-section $\times$ BR in Table 1 for different masses of the scalars. Some of the processes also get limits from ATLAS study [ATLAS:2021jol]. We have shown these limits in parentheses, in Table 1, comparing processes with the same initial and final states, in our model and the ATLAS study. We cite the limits from Ref. [ATLAS:2021jol] in $4l$, $3l+\geq 2j$ and $2l+\geq 3j$ channels. Note that the limit in $2l+\geq 3j$ channel is applicable only when the two leptons have the same sign. We also did not include any channel with jet multiplicity greater than 4 because these channels suffer from large SM background.
@@
Table 1: fermiophobic scenario: Cross-section $\times$ BR for various final states at the $14$ TeV LHC, for $M_{\Phi}=300,~500$ and $800$ GeV, with $M_{\Sigma}=M_{\Phi}-80$ GeV. We use $y_{\nu}=0.0001$. The values in parentheses indicate the existing upper limits from ATLAS [ATLAS:2021jol] .
We find that the pair production cross-section of $\phi_{1}^{+}\phi_{1}^{-}$ does not receive any limit from the LHC. The doubly charged scalar with mass 300 GeV is excluded from the ATLAS search in all three channels when produced in pair. However, doubly charged scalars with mass 500 GeV are not excluded in these channels. In the associated production mode also, the 300 GeV mass is excluded in $3l+4j$, $3l+2j$ and $2l_{SC}+4j$ channels from $3l~+~\geq 2j$ and $2l~+~\geq 3j$ channel searches of ATLAS. The multilepton channels do not face any constraints from the LHC, but the cross-sections in these channels (with four or more leptons) remain significantly small.
There exists no limit from the LHC on the remaining channels, involving four or more leptons. Note that these final state leptons originate from $W/Z$ boson decays and thus, these channels experience suppression due to small branching ratio. Only the channel with four leptons and two jets offers a better cross-section $\times$ BR. However, as two of the leptons come from the decay of $W$ boson, the reconstruction of the quartet scalar or the quintuplet fermion is not straightforward. In the fermiophilic scenario, which we discuss next, the leptons are produced directly from the decay of the scalar or fermion, resulting in a large cross-section $\times$ BR. This also facilitates the reconstruction of the BSM particles.
@@
The other possibilities are $pp\rightarrow\phi_{1}^{+}\phi_{1}^{-}\rightarrow\Sigma^{+}\nu~\Sigma^{-}\overline{\nu}$ and $pp\rightarrow\phi_{1}^{+}\phi_{1}^{-}\rightarrow\Sigma^{0}l^{+}~\Sigma^{-}\overline{\nu}$. We have observed in the previous section that the branching ratio for $\phi_{1}^{+}\rightarrow\Sigma^{+}\nu$ is larger than $\phi_{1}^{+}\rightarrow\Sigma^{0}l^{+}$. The branching ratio for $\Sigma^{\pm}$ decays to $W^{\pm}\nu(\overline{\nu})$ is almost 95%. Hence, the final states from these decay modes will be $W^{+}W^{-}$+ MET or $W^{\pm}W^{-}l^{\mp}l^{+}$ + MET. There will be large MET in final states if the $W$s decay leptonically and reconstruction of the scalar mass will not be straightforward. On the other hand, $W$s decaying to the jets will lead to final states which suffer from large SM backgrounds.
The other possible fermiophilic channels are,
$pp\rightarrow\phi^{++}\phi^{--}$ $\rightarrow$ $\Sigma^{++}\nu~\Sigma^{--}\overline{\nu}~~\rightarrow W^{+}l^{+}\nu W^{-}l^{-}\overline{\nu}$ (24)
$pp\rightarrow\phi^{++}\phi_{2}^{-}$ $\rightarrow$ $\Sigma^{++}\nu~\Sigma^{--}l^{+}\rightarrow W^{+}l^{+}\nu W^{-}l^{-}l^{+}$ (25)
The cross-section $\times$ BR in different channels are presented in Table 2. We do not consider the pair production of $\phi_{2}^{+}\phi_{2}^{-}$ because the cross-section is one order of magnitude smaller. Some of the processes also get limit from the ATLAS study [ATLAS:2021jol] which are shown in the parentheses. However, these limits are only meant for a preliminary comparison. Even though the final states are same, the exact signal topology, particularly in the fermiophilic scenario, is different from that in the ATLAS study.
We observe from [ATLAS:2021jol] that the 1st process (Eq. 23) does not get any limits from the LHC but the cross-section $\times$ BR is low compared to the other channels. The doubly charged scalars with 500 GeV are excluded by the ATLAS search in $3l+\geq 2j$ and $4l$ channels  [^8] ATLAS limit does not apply to $2l+4j$ state here because ATLAS considers only same charge leptons, whereas in our study we have two oppositely charged leptons.. The $2l+4j$ channel with two opposite sign leptons suffer from a very high SM background. Also, in the associated production mode, scalars with 300 GeV mass are excluded in the $3l+4j$ channel.
@@
Table 2: fermiophilic scenario: Cross-section $\times$ BR for various final states at the $14$ TeV LHC, for $M_{\Phi}=300,~500$ and $800$ GeV, with $M_{\Sigma}=M_{\Phi}-80$ GeV. We use parameter values $y_{\nu}=1$ and $v_{4}=0.0002$. The values in parentheses indicate the existing upper limits from ATLAS [ATLAS:2021jol].
The multilepton final states with 4 to 5 leptons and 2 or more jets are not excluded by the LHC searches but they suffer from a very small branching ratio. Note that, in the fermiophilic signals, all or most of the SM leptons originate from the decay of the scalars or the quintuplet fermions while the jets come from the decays of $W/Z$ bosons. Hence, the branching ratio is larger compared to the fermiophobic scenario (Table 2). The mass of the charged scalars and the quintuplet fermions can be reconstructed easily in the fermiophilic scenario.
@@ 4 Analysis and Results
The pair production of $\phi_{1}^{+}\phi_{1}^{-}$ and associated production $\phi^{++}\phi_{2}^{-}$ is shown in Fig. 3 and Fig. 4 for the fermiophobic and fermiophilic scenarios, respectively. In this section, we discuss the detailed collider analysis corresponding to the fermiophilic scenario as depicted in Fig. 4. We consider pair-production of $\phi_{1}^{\pm}$ as our first signal process (S1) where $\phi_{1}$ undergoes a cascade decay through the intermediate state $\Sigma^{0}$, resulting in a final state with four charged leptons and two W bosons (Fig. 4(a)). In addition to S1, we also study the associated production of $\phi_{2}^{\pm}$ with $\phi^{\mp\mp}$, which constitutes our second signal process (S2), resulting in a final state with three charged leptons and two W bosons. Note that we have also considered the charged conjugated process as well.
@@
The processes in Fig. 4 results in a final state containing four/five leptons and two jets when one of the two $W$ bosons decay hadronically while the other decays leptonically. For S1, both the possibilities are considered. There is no other source of MET in S1 apart from the decay of leptonically decaying $W$. $W$ can be reconstructed from the jets and that can be combined with correct leptons to reconstruct the fermion or scalar. We will discuss the reconstruction later.
@@
Since both signal processes produce final states with at least four leptons accompanied by jets and missing transverse energy (MET), the most relevant SM backgrounds are those that can mimic such signatures. Therefore, we consider SM processes containing at least four leptons and two jets, along with a non-zero source of MET, as potential backgrounds to S1 and S2. The list of dominant SM backgrounds includes $t\overline{t}Z$, $tWZj$, ZWWjjZWWjj, ZZWjjZZWjj, $t\overline{t}WW$, and $t\overline{t}t\overline{t}$, which exhibit similar final-state topologies with cross-sections  [ATLAS:2023dkz, CMS:2020fqz] comparable to (in some cases larger) the signal S1 and S2. A summary of the signal and background processes, along with their relevant decay modes and cross-section $\times$ BR, is provided in Table 3.
@@
Table 3: The cross-sections $(\sigma\times$ BR) for the two signal (S1 and S2) processes are presented for three benchmark points. Background processes are listed as (B1-B6) and their cross-sections $(\sigma\times$ BR) are also given.
@@
The event samples for both signals and backgrounds are generated (in leading-order in $\alpha_{s}$) using MadGraph (version 3.4.1)[Alwall:2011uj] at a center-of-mass energy of 14 TeV. The effects of parton showering, hadronization, as well as initial state radiation (ISR) and final state radiation (FSR) are simulated using PYTHIA8 (version 8.310)[Bierlich:2022pfr]. The detector-level effects are incorporated using Delphes (version 3.5.0)[deFavereau:2013fsa], employing the default CMS detector configuration. The jets are reconstructed using the anti-$k_{T}$ algorithm as implemented in FastJet [Cacciari:2005hq, Cacciari:2011ma] with a radius parameter $R=0.5$ and a minimum transverse momentum for the jets set at $p^{\text{min}}_{T,j}=20~\text{GeV}$. All events are required to pass through a few basic selection cuts for the leptons and jets, as: $p_{T,l}>10$ GeV, $|\eta_{l}|<2.5$ and $p_{T,j}>30$ GeV, $|\eta_{j}|<4.5$, where $p_{T,l}$ , $p_{T,j}$ are the transverse momentum of the leptons and jets, and $\eta_{l}$, $\eta_{j}$ are their pseudorapidities, respectively. The jets are tagged as b-jets using the default b-tagging module of Delphes CMS card.
@@
In our analysis, leptons play an important role, and therefore a proper identification and selection strategy for the leptons is extremely crucial. Given the mass difference between the scalar and the fermion, the use of the MET direction in identifying the leptons turned out to be an efficient methodology. We therefore arrange the leptons in an increasing order of magnitude of $\Delta\phi_{l,MET}$, considering that the lepton with the smallest separation in $\Delta\phi_{l,MET}$ is produced by the decay W-boson. To further iterate this point, let us consider the process S2, which has two sources of missing energy, one from the decay of $\phi^{++}$ and another from W-boson decay. But the mass difference between $\phi^{++}\;\rm and\;\Sigma^{++}$ is smaller than that of $\Sigma^{++}$ and the W-boson. Hence it is expected that the lepton from the decay of $W$ will be identified in the event as the $0^{th}$ lepton ($l_{0}$)following the above mentioned principle of numbering the leptons. Hence, we obtain 4 leptons as ($l_{0}$$l_{1}$$l_{2}$$l_{3}$, $l_{4}$), where $l_{0}$ is the closest to the MET direction or in other words, furthest from the di-jet branch, where W decays to jets.
In the upper panel of Fig. 5, we show the normalized distribution of the multiplicity of leptons ($N_{lep}$), jets ($N_{jet}$), and b-tagged jets ($N_{b-jet}$). The signal S1 is plotted for two reference points of scalar mass 300 and 500 GeV, while S2 for the mass of 500 GeV. For the backgrounds, only the first three dominant ones are considered in these plots, which helps to identify the appropriate cuts to be applied for event selection and in the optimization process. Several key points to note from these three distributions:
@@
(iii) As discussed before, the invariant mass of the leptons could play an important role in isolating signal events from the background ones. Therefore, we plot the invariant mass of the leading (ordered with respect to the direction with MET) two opposite sign lepton pairs ($M_{l^{+}l^{-}}$), in the lower panel of Fig. 5. Most of the backgrounds, B1, B2, B3 and B4, involve a Z-boson which can be reconstructed using opposite-sign leptons. Hence, we observe a clear resonance peak at 90 GeV for the backgrounds, unlike the signal events, where the leptons are not coming from the decay of the $Z$ boson.
@@
Figure 5: Normalized distributions of the variables used in the pre-selection: Lepton multiplicity (Top left), jet multiplicity (Top middle), b-jet multiplicity (Top right), the invariant mass of the two opposite-sign lepton pairs, (Bottom left and Bottom right).
@@
  • Angular separation between two jets ($\Delta R_{jj}$): Signal events (S1 and S2) produce jets from the decay of W-boson that originated from a heavy fermion, and therefore, these jets are expected to be collimated. As expected, we observe a peak around $\Delta R_{jj}=0.5$ for signal events with a sharply falling distribution after the peak (see the lower middle figure of Fig. 6). The jets from the background events have a relatively flat distribution in $\Delta R_{jj}$. This characteristic difference in the distribution of $\Delta R_{jj}$ makes it an important variable to distinguish signal events from background events.
  • Invariant mass of a pair of jets ($M_{jj}$): Following the argument given above, the invariant mass of the leading pair of jets (having smallest $\Delta R$ between them) is also expected to peak at the W-boson mass. The peak is clearly visible for the signal events in the distribution shown in the top right panel of Fig. 7, unlike the background events.
  • $p_{T}$ of the leptons: As we have discussed before, we had identified the leptons as $l_{0}$$l_{1}$$l_{2}$$l_{3}$ and $l_{4}$, where $l_{0}$ is the closest to the MET direction or in other words, furthest from the di-jet branch, where W decays to jets. We further arrange the leptons in the increasing value of $p_{T}$ and we show the distributions in Figure 6. However we do not use a different notation, we keep the previous notation for the ordered leptons.
  • Invariant mass of a pair of leptons ($M_{ll}$):
    The same-sign leptons do not originate from any particular decay chain. In the signal processes, same-sign leptons can arise from the decays of exotic particles or W, whereas for the backgrounds, same-sign leptons originate predominantly from Standard Model $W$ and $Z$ boson decays. Figure 7 (top right panel) shows the invariant mass distribution of the two leading same-sign leptons (ordered according to their direction with respect to the missing transverse energy (MET)), for both the signal and background processes. The pattern of the invariant mass distributions exhibit a significant distinction between signal and background topologies. As the scalar mass increases, the signal distribution shifts towards higher masses relative to the background. Furthermore, for the same benchmark mass of 500 GeV, the distribution for S2 is more separated from the background distributions compared to S1.
@@
Understanding the signals at different benchmark points lead us to identify the features useful to discriminate the signal from the backgrounds having similar final states but higher cross-sections. Now we focus on building a classification model, which can help the signals achieve an impressive significance even if the cross-section is small. The combined data set of a signal for a specific benchmark point together with the backgrounds are divided into a $50:50$ ratio for training and testing in such a way that for training there are equal numbers of signal and background events present. Each of the background is weighted according to their cross-section with the weight = $\frac{1}{N_{MC}}\times\sigma_{MC}\times\mathcal{BR}\times(\int\mathcal{L}dt)$, where $N_{MC}$ is the number of Monte Carlo samples generated, $\sigma_{MC}$ is the production cross-section of each process, $\mathcal{BR}$ is the branching ratio for a particular decay channel and $\int\mathcal{L}dt$ is the integrated luminosity. The analysis is carried out on the high luminosity run of LHC with $\int\mathcal{L}dt\;=\;3000fb^{-1}$. All features or observables are passed through the MVA, and a Receiver Operating Characteristics (ROC) curve is obtained that gives the probability of correctly rejecting background events as a function of the probability of correctly identifying signal events. We use the ROOT Library, “Toolkit for Multivariate Analysis” [TMVA:2007ngy] to perform the MVA. More specifically, we use the algorithm “Adaptive Boosted Decision Tree” (AdaBoost) [FREUND1997119] for the classification of signal and background events. In Fig. 8, we show the ROCs obtained for the signal process S1 and S2 for two representative values of the scalar mass. As evident from the figure, the higher the scalar mass, the better the separation in the input observables, leading to improved performances in terms of the Area Under the Curve (AUC).
To estimate the discovery potential of the HL-LHC, we calculate the signal significances for each benchmark point, defined as (also referred to as the Asimov significance [Cowan:2010js]):

### v2

@@ 3.3 Existing Collider Searches
In the model under inspection, the scalars decay either to the gauge bosons directly or to the quintuplet fermions which subsequently decay to gauge bosons, as we have discussed in the previous section. Before discussing possible signatures at the LHC, let us discuss the existing limits from collider experiments on charged scalars. The direct search limit on charged Higgs comes from LEP which is around 80 GeV [13]. Both ATLAS and CMS have conducted extensive searches targeting multiple production and decay modes of the charged scalars. A recent summary of these searches is provided in [68, 12, 47, 20]. Most LHC searches target single production of the charged scalars via vector-boson fusion and have put limits in cases where the scalar decays to gauge bosons[9, 11, 67, 3]. In this case, a lower limit of $\sim 200$ GeV is placed on the masses of charged scalars. Limits also exist on production cross sections of charged scalars when they directly decay to SM fermions [65, 4, 2, 1, 6, 8] or specifically, to quarks [5]. These searches have placed a lower limit of $\sim 400$ GeV on the charged scalar masses.
Ref. [7] has studied the doubly charged scalar pair production as well as associated production of singly and doubly charged scalars, with each scalar decaying to gauge bosons and limits are placed from the searches in 4 leptons, 3 leptons + $\geq$ 2 jets and 2 leptons $\geq$ 3 jets final states. As no excess has been observed, limits have been placed on the cross section $\times$ BR (See Fig. 9 of [7]).
@@ 3.4 Fermiophobic and Fermiophilic Signatures
Let us consider the fermiophobic scenario when $M_{\Phi}>M_{\Sigma}$ and the coupling of the scalars with the quintuplet fermion is small, $y_{\nu}<0.005$. In this situation, the branching ratio of the scalars to the SM fermions are negligible. This situation gives rise to the final states with gauge bosons
$pp~\rightarrow~\phi_{1}^{+}\phi_{1}^{-}$ $\rightarrow$ $W^{+}Z~W^{-}Z,$ (20)
$pp\rightarrow\phi^{++}\phi^{--}$ $\rightarrow$ $W^{+}W^{+}W^{-}W^{-},$ (21)
@@
We present the possible final states and the values of cross section $\times$ BR in Table 1 for different masses of the scalars. Some of the processes also get limits from ATLAS study [7]. We have shown these limits in parentheses, in Table 1, comparing processes with the same initial and final states, in our model and the ATLAS study. We cite the limits from Ref. [7] in $4l$, $3l+\geq 2j$ and $2l+\geq 3j$ channels. Note that the limit in $2l+\geq 3j$ channel is applicable only when the two leptons have the same sign. We also did not include any channel with jet multiplicity greater than 4 because these channels suffer from large SM background.
@@
Table 1: fermiophobic scenario: Cross section $\times$ BR for various final states at the $14$ TeV LHC, for $M_{\Phi}=300,~500$ and $800$ GeV, with $M_{\Sigma}=M_{\Phi}-80$ GeV. We use $y_{\nu}=0.0001$. The values in parentheses indicate the existing upper limits from ATLAS [7] .
We find that the pair production cross section of $\phi_{1}^{+}\phi_{1}^{-}$ does not receive any limit from the LHC. The doubly charged scalar with mass 300 GeV is excluded from the ATLAS search in all three channels when produced in pair. However, doubly charged scalars with mass 500 GeV are not excluded in these channels. In the associated production mode also, the 300 GeV mass is excluded in $3l+4j$, $3l+2j$ and $2l_{SC}+4j$ channels from $3l~+~\geq 2j$ and $2l~+~\geq 3j$ channel searches of ATLAS. The multilepton channels do not face any constraints from the LHC, but the cross sections in these channels (with four or more leptons) remain significantly small.
There exists no limit from the LHC on the remaining channels, involving four or more leptons. Note that these final state leptons originate from $W/Z$ boson decays and thus, these channels experience suppression due to small branching ratio. Only the channel with four leptons and two jets offers a better cross section $\times$ BR. However, as two of the leptons come from the decay of $W$ boson, the reconstruction of the quartet scalar or the quintuplet fermion is not straightforward. In the fermiophilic scenario, which we discuss next, the leptons are produced directly from the decay of the scalar or fermion, resulting in a large cross section $\times$ BR. This also facilitates the reconstruction of the BSM particles.
@@
The other possibilities are $pp\rightarrow\phi_{1}^{+}\phi_{1}^{-}\rightarrow\Sigma^{+}\nu~\Sigma^{-}\overline{\nu}$ and $pp\rightarrow\phi_{1}^{+}\phi_{1}^{-}\rightarrow\Sigma^{0}l^{+}~\Sigma^{-}\overline{\nu}$. We have observed in the previous section that the branching ratio for $\phi_{1}^{+}\rightarrow\Sigma^{+}\nu$ is larger than $\phi_{1}^{+}\rightarrow\Sigma^{0}l^{+}$. The branching ratio for $\Sigma^{\pm}$ decays to $W^{\pm}\nu(\overline{\nu})$ is almost 95%. Hence, the final states from these decay modes will be $W^{+}W^{-}$+ MET or $W^{\pm}W^{-}l^{\mp}l^{+}$ + MET. There will be a large MET in the final states if the $W$s decay leptonically, and thereby, reconstruction of the scalar mass will not be straightforward. On the other hand, $W$s decaying to the jets will lead to final states, which suffer from large SM backgrounds.
The other possible fermiophilic channels are
$pp\rightarrow\phi^{++}\phi^{--}$ $\rightarrow$ $\Sigma^{++}\nu~\Sigma^{--}\overline{\nu}~~\rightarrow W^{+}l^{+}\nu W^{-}l^{-}\overline{\nu},$ (24)
$pp\rightarrow\phi^{++}\phi_{2}^{-}$ $\rightarrow$ $\Sigma^{++}\nu~\Sigma^{--}l^{+}\rightarrow W^{+}l^{+}\nu W^{-}l^{-}l^{+}.$ (25)
The cross section $\times$ BR in different channels are presented in Table 2. We do not consider the pair production of $\phi_{2}^{+}\phi_{2}^{-}$ because the cross section is one order of magnitude smaller. Some of the processes also get a limit from the ATLAS study [7], which are shown in parentheses. However, these limits are only meant for a preliminary comparison. Even though the final states are the same, the exact signal topology, particularly in the fermiophilic scenario, is different from that in the ATLAS study.
We observe from [7] that the 1st process (Eq. 23) does not get any limits from the LHC but the cross section $\times$ BR is low compared to the other channels. The doubly charged scalars with 500 GeV are excluded by the ATLAS search in $3l+\geq 2j$ and $4l$ channels  [^8] ATLAS limit does not apply to $2l+4j$ state here because ATLAS considers only same charge leptons, whereas in our study we have two oppositely charged leptons.. The $2l+4j$ channel with two opposite sign leptons suffers from a very high SM background. Also, in the associated production mode, scalars with 300 GeV mass are excluded in the $3l+4j$ channel.
@@
Table 2: fermiophilic scenario: Cross section $\times$ BR for various final states at the $14$ TeV LHC, for $M_{\Phi}=300,~500$ and $800$ GeV, with $M_{\Sigma}=M_{\Phi}-80$ GeV. We use parameter values $y_{\nu}=1$ and $v_{4}=0.0002$. The values in parentheses indicate the existing upper limits from ATLAS [7].
The multilepton final states with 4 to 5 leptons and 2 or more jets are not excluded by the LHC searches, but they suffer from a very small branching ratio. Note that, in the fermiophilic signals, all or most of the SM leptons originate from the decay of the scalars or the quintuplet fermions, while the jets come from the decays of $W/Z$ bosons. Hence, the branching ratio is larger compared to the fermiophobic scenario (Table 2). The mass of the charged scalars and the quintuplet fermions can be reconstructed easily in the fermiophilic scenario.
@@ 4 Analysis and Results
The pair production of $\phi_{1}^{+}\phi_{1}^{-}$ and associated production $\phi^{++}\phi_{2}^{-}$ is shown in Fig. 3 and Fig. 4 for the fermiophobic and fermiophilic scenarios, respectively. In this section, we discuss the detailed collider analysis corresponding to the fermiophilic scenario as depicted in Fig. 4. We consider pair production of $\phi_{1}^{\pm}$ as our first signal process (S1) where $\phi_{1}$ undergoes a cascade decay through the intermediate state $\Sigma^{0}$, resulting in a final state with four charged leptons and two W bosons (Fig. 4(a)). In addition to S1, we also study the associated production of $\phi_{2}^{\pm}$ with $\phi^{\mp\mp}$, which constitutes our second signal process (S2), resulting in a final state with three charged leptons and two W bosons. Note that, we have considered the charged conjugate process as well.
@@
The processes in Fig. 4 results in a final state containing four/five leptons and two jets when one of the two $W$ bosons decays hadronically while the other decays leptonically. For S1, both the possibilities are considered. There is no other source of MET in S1 apart from the decay of leptonically decaying $W$. $W$ can be reconstructed from the jets, and can be combined with correct leptons to reconstruct the fermion or scalar. We will discuss the reconstruction later.
@@
Since both signal processes produce final states with at least four leptons accompanied by jets and missing transverse energy (MET), the most relevant SM backgrounds are those that can mimic such signatures. Therefore, we consider SM processes containing at least four leptons and two jets, along with a non-zero source of MET, as potential backgrounds to S1 and S2. The list of dominant SM backgrounds includes $t\overline{t}Z$, $tWZj$, ZWWjjZWWjj, ZZWjjZZWjj, $t\overline{t}WW$, and $t\overline{t}t\overline{t}$, which exhibit similar final-state topologies with cross sections  [10, 66] comparable to (in some cases larger) the signal S1 and S2. A summary of the signal and background processes, along with their relevant decay modes and cross section $\times$ BR, is provided in Table 3.
@@
Table 3: The cross sections $(\sigma\times$ BR) for the two signal (S1 and S2) processes are presented for three benchmark points. Background processes are listed as (B1-B6) and their cross sections $(\sigma\times$ BR) are also given.
@@
The event samples for both signals and backgrounds are generated (in leading-order in $\alpha_{s}$) using MadGraph (version 3.4.1)[16] at a center-of-mass energy of 14 TeV. The effects of parton showering, hadronization, as well as initial state radiation (ISR) and final state radiation (FSR) are simulated using PYTHIA8 (version 8.310)[22]. The detector-level effects are incorporated using Delphes (version 3.5.0)[32], employing the default CMS detector configuration. The jets are reconstructed using the anti-$k_{T}$ algorithm as implemented in FastJet [27, 26] with a radius parameter $R=0.5$ and a minimum transverse momentum for the jets set at $p^{\text{min}}_{T,j}=20~\text{GeV}$. All events are required to pass through a few basic selection cuts for the leptons and jets, as: $p_{T,l}>10$ GeV, $|\eta_{l}|<2.5$ and $p_{T,j}>30$ GeV, $|\eta_{j}|<4.5$, where $p_{T,l}$ , $p_{T,j}$ are the transverse momentum of the leptons and jets, and $\eta_{l}$, $\eta_{j}$ are their pseudorapidities, respectively. The jets are tagged as b-jets using the default b-tagging module of Delphes CMS card.
@@
In our analysis, leptons play an important role and therefore, a proper identification and selection strategy for the leptons is extremely crucial. Given the mass difference between the scalar and the fermion, the use of the MET direction in identifying the leptons turned out to be an efficient methodology. We, therefore, arrange the leptons in an increasing order of magnitude of $\Delta\phi_{l,MET}$, considering that the lepton with the smallest separation in $\Delta\phi_{l,MET}$ is produced by the decay of W-boson. To further iterate this point, let us consider the process S2, which has two sources of missing energy; one from the decay of $\phi^{++}$ and another from W-boson decay. However, the mass difference between $\phi^{++}\;\rm and\;\Sigma^{++}$ is smaller than that of $\Sigma^{++}$ and the W-boson. Hence, it is expected that the lepton from the decay of $W$ will be identified in the event as the $0^{th}$ lepton ($l_{0}$), following the above mentioned rule of numbering the leptons. Hence, we obtain 4 leptons as ($l_{0}$$l_{1}$$l_{2}$$l_{3}$, $l_{4}$), where $l_{0}$ is the closest to the MET direction, or in other words, farthest from the di-jet branch, where W decays to jets.
In the upper panel of Fig. 5, we show the normalized distribution of the multiplicity of leptons ($N_{lep}$), jets ($N_{jet}$), and b-tagged jets ($N_{b-jet}$). The signal S1 is plotted for two reference points of scalar mass viz., 300 and 500 GeV, while S2 is plotted for the mass 500 GeV. For the backgrounds, only the first three dominant ones are considered in these plots, which helps to identify the appropriate cuts to be applied for event selection and in the optimization process. Several key points to note from these three distributions:
@@
(iii) As discussed before, the invariant mass of the leptons could play an important role in isolating signal events from the background ones. Therefore, we plot the invariant mass of the leading (ordered with respect to the direction with MET) two opposite-sign lepton pairs ($M_{l^{+}l^{-}}$), in the lower panel of Fig. 5. Most of the backgrounds, B1, B2, B3 and B4, involve a Z-boson which can be reconstructed using opposite-sign leptons. Hence, we observe a clear resonance peak at 90 GeV for the backgrounds, unlike the signal events, where the leptons are not coming from the decay of the $Z$ boson.
@@
Figure 5: Normalized distributions of the variables used in the pre-selection: Lepton multiplicity (Top left), jet multiplicity (Top middle), b-jet multiplicity (Top right ), the invariant mass of the two opposite-sign lepton pairs (Bottom left and Bottom right).
@@
  • Angular separation between two jets ($\Delta R_{jj}$):
    Signal events (S1 and S2) produce jets from the decay of a W boson that originated from a heavy fermion, and therefore, these jets are expected to be collimated. As expected, we observe a peak around $\Delta R_{jj}=0.5$ for signal events with a sharply falling distribution after the peak (see the lower middle figure of Fig. 6). The jets from the background events have a relatively flat distribution in $\Delta R_{jj}$. This characteristic difference in the distribution of $\Delta R_{jj}$ makes it an important variable to distinguish signal events from background events.
  • Invariant mass of a pair of jets ($M_{jj}$): Following the argument given above, the invariant mass of the leading pair of jets (having the smallest $\Delta R$ between them) is also expected to peak at the W-boson mass. The peak is clearly visible for the signal events in the distribution shown in the top right panel of Fig. 7, unlike the background events.
  • $p_{T}$ of the leptons: As discussed before, we identify the leptons as $l_{0}$, $l_{1}$, $l_{2}$, $l_{3}$, and $l_{4}$, where $l_{0}$ is the closest to the MET direction, or in other words, farthest from the di-jet branch, where W decays to jets. We further arrange the leptons in the increasing value of $p_{T}$ and we show the distributions in Figure 6. However, we do not use a different notation, we keep the previous notation for the ordered leptons.
  • Invariant mass of a pair of leptons ($M_{ll}$): The same-sign leptons do not originate from any particular decay chain. In the signal processes, same-sign leptons can arise from the decays of exotic particles or W, whereas for the backgrounds, same-sign leptons originate predominantly from Standard Model $W$ and $Z$ boson decays. Figure 7 (top right panel) shows the invariant mass distribution of the two leading same-sign leptons (ordered according to their direction with respect to the missing transverse energy (MET)), for both the signal and background processes. The pattern of the invariant mass distributions exhibits a significant distinction between signal and background topologies. As the scalar mass increases, the signal distribution shifts towards higher masses relative to the background. Furthermore, for the same benchmark mass of 500 GeV, the distribution for S2 is more separated from the background distributions compared to S1.
@@
Understanding the signals at different benchmark points leads us to identify the features useful for discriminating the signal from the backgrounds having similar final states but higher cross sections. Now, we focus on building a classification model, which can help the signals achieve an impressive significance even if the cross section is small. The combined data set of a signal for a specific benchmark point, together with the backgrounds, are divided into a $50:50$ ratio for training and testing in such a way that for training, there are equal numbers of signal and background events present. Each of the background is weighted according to their cross section with the weight = $\frac{1}{N_{MC}}\times\sigma_{MC}\times\mathcal{BR}\times(\int\mathcal{L}dt)$, where $N_{MC}$ is the number of Monte Carlo samples generated, $\sigma_{MC}$ is the production cross section of each process, $\mathcal{BR}$ is the branching ratio for a particular decay channel and $\int\mathcal{L}dt$ is the integrated luminosity. The analysis is carried out on the high luminosity run of LHC with $\int\mathcal{L}dt\;=\;3000fb^{-1}$. All features or observables are passed through the MVA, and a Receiver Operating Characteristics (ROC) curve is obtained that gives the probability of correctly rejecting background events as a function of the probability of correctly identifying signal events. We use the ROOT Library, “Toolkit for Multivariate Analysis” [46] to perform the MVA. More specifically, we use the algorithm “Adaptive Boosted Decision Tree” (AdaBoost) [36] for the classification of signal and background events. In Fig. 8, we show the ROCs obtained for the signal process S1 and S2 for two representative values of the scalar mass. As evident from the figure, the higher the scalar mass, the better the separation in the input observables, leading to improved performances in terms of the Area Under the Curve (AUC).
To estimate the discovery potential of the HL-LHC, we calculate the signal significances for each benchmark point, defined as (also referred to as the Asimov significance [31]):